OpenStudy (anonymous):

What is the spring constant for the spring in the graph? Will give medal! See my reply post for graph!

OpenStudy (anonymous):

See attached file for graph.

OpenStudy (anonymous):

im not sure how to do this but I will try to get someone over here to help you :))

OpenStudy (anonymous):

@myininaya @AllTehMaffs @amistre64

OpenStudy (anonymous):

So, if this were just a line, what would its basic equation of a line look like  ?

OpenStudy (anonymous):

umm I'm not really sure but I think it would be the slope and for the question I asked I think it would be 25.0 N/m not sure though

OpenStudy (anonymous):

or +0.04 N/m

OpenStudy (anonymous):

^_^ The equation of this line would fit \[ y=mx+b\] So looking at your graph, just replace these variables with force for y and distance for x giving you \[F=-kx\] which is Hooke's law. And you're correct in that the slope would be the slope of the line. \[ -k=\frac{\Delta y}{\Delta x}=\frac{-.5N}{.02m}\] \[k=25 \ N/m\]

OpenStudy (anonymous):

~the spring constant would be the slope of the line

OpenStudy (anonymous):

oh ok got it thanks!

OpenStudy (anonymous):

Welcome ^_^

OpenStudy (anonymous):

can u help me with one more question?

OpenStudy (anonymous):

yah yah

OpenStudy (anonymous):

What is the period of the spring's motion in the position-time graph shown here?

OpenStudy (anonymous):

is it 2.0 s?

OpenStudy (anonymous):

The period of a function like that is from one crest to the next